The Rich Corporation has $150 million worth of common stock on which investors require a 15% rate of return. It has $35 million in bonds that offer a 5% return.

a.

Calculate the WACC for Rich Corporation if they are subject to a 35% tax rate.

b.

Recompute Richs' WACC assuming the firm has $95 million in debt and $90 million in stock.

c.

After reviewing parts a. and b. explain why the WACC calculated in part b. may not be the correct answer if the capital structure changes.
Answer #1

(a) Common Stock = E = $150 m

Debt = D = $35 million

we = weight firm’s equity = 150 / (150+35) = 0.8108

wd = weight the firm’s debt = 35 / (150+35) = 0.1892

Re = cost of equity = 0.15

Rd = cost of debt = 0.05

Tax Rate = T = 35%

WACC = we*Re + wd*Rd*(1-T)

=> WACC = 0.8108*0.15 + 0.1892*0.05*(1-0.35) = 0.1278 = 12.78%

(b)

Common Stock = E = $90 m

Debt = D = $95 million

we = weight firm’s equity = 90 / (90+95) = 0.4865

wd = weight the firm’s debt = 95 / (90+95) = 0.5135

Re = cost of equity = 0.15

Rd = cost of debt = 0.05

Tax Rate = T = 35%

WACC = we*Re + wd*Rd*(1-T)

=> WACC = 0.4865*0.15 + 0.5135*0.05*(1-0.35) = 0.0897 = 8.97%

(c) When WACC is calculated for a firm, it is assumed that the company's current mix of debt and equity capital will remain same in the future. We are changing the capital structure in part (b) and hence the WACC calculated in part (b) may not be correct.